Fermi National Laboratory
 

Accelerator Update for January 31, 2007 - February 2, 2007

 

Wednesday January 31

The day shift began with Operations monitoring the Tevatron (TeV) store 5201 (40.88E30), with the Antiproton Source (Pbar) stacking (139.8E10), with the Main Injector (MI), Recycler, and NuMI off for HV101 repair, with the Linac and Booster operating normally, with MiniBooNE
off, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

Operators prepared the Meson enclosures M03, 4, & 5 for access at 9:47 AM.

At 11:26 AM, a TeV RF station (TRF1) tripped off and caused store 5101 to abort. There was no quench. (TRF8 had already tripped off earlier.)

Operations began preparing the TeV for access at 12:29 PM.

Technicians and experts accessed the TeV at 12:44 PM to complete the sextupole circuit wiring.

At 2:41 PM, Operations reported a surge in helium and vacuum problems in TeV sector A1. Cryo system technicians reported that they could not keep the system up with the poor vacuum.

At 2:59 PM, Cryo system techs reported that they were turning off and beginning the warm up of the TeV sector A1 cryo system.

At 3:03 PM, a TeV expert reported a vacuum leak on from a vacuum pump in sector A18.

At 4:42 PM, two Booster RF stations (BRF7 & 8) tripped off due to a vacuum pump (IPRF17) trip. Everything reset okay.

At 5:16 PM, the H- Source output went to zero. Operators called in experts.

Preaccelerator experts had the H- source back up and beam to the Booster absorber by 7 PM. They found and repaired a bad high voltage cable on the extractor power supply. Unfortunately, the output of the source came back at a lower intensity than normal. An expert tuned it to a stable output of ~45 mA.

Techs reported that the HV101 magnet electrical load cable connections had been completed at 6:28 PM.

Operations reported at 8:06 PM that TeV sector A2 cryo system had been warmed up to 100 Kelvin. Cryo system techs at CHL will next warm up sectors A1 and A3.

At 10:24 PM, Operations reported that the vacuum for the sector A17 separators was going bad. Operations had been conditioning the separators. Operators turned off the separators and called in an expert.

Thursday February 1

The midnight shift began with Operations monitoring the Pbar stack (133.2E10), with TeV sectors A1 and 3 warming up, and with HV101 repairs continuing.

Operations reported at 6:34 AM that all the shielding except two top pieces had been installed to HV101. The vacuum pumps were on and pumping the area down.

At 10:42 AM, a Linac quadrupole power supply’s (QPSC02) current went low and caused the loss of the Neutron Therapy Facility's (NTF) beam. Operations adjusted the current.

At 11:51 AM, NTF’s QPSC02 current again drifted low. Operators contacted an expert who instructed them to lower the power supply’s setting.

Operations reported at 12:01 PM that the TeV vacuum pump (A1IP8) in sector A1 that was leaking had been replaced.

NTF reported at 12:31 PM that they had completed their last patient treatment.

Operators had completed a search and secure of NuMI, MI, and the Pelletron by 4:30 PM.

Operations established beam to the MI absorber at 5:33 PM.

Operations reported at 6:10 PM that Recycler experts were using protons for studies.

Operations reported at 6:12 PM that Pbar was stacking.

At 7:11 PM, a NuMI expert reported a problem with a beam monitor in the MI tunnel that required an access. Operations halted the MI beam at 7:39 PM and prepared the area for an access.

Operations reported at 9:22 PM that the NuMI work had been completed. Operators established a MI ramp.

Operations established beam to Pbar and protons to the Recycler by 9:30 PM.

Operations established beam to NuMI at 9:42 PM.

The Recycler studies finished at 11:29 PM. Operations began preparing the Recycler to stash antiprotons.

Friday February 2

The midnight shift began with the TeV off, with Operations monitoring the vacuum at TeV sector A1 and the stash (108.8E10), with Pbar stacking (9.3E10), with MI, Recycler, Linac, and Booster operating normally, with NuMI taking beam, with MiniBooNE off, and with CDF and D0 in access.

At 4:01 AM, a Booster RF station (BRF6) tripped off. It reset okay.

At 5:13 AM, the West Anode Power Supply tripped off and wouldn’t reset. Operators investigated and found that BRF9 was off on a ground fault indication. They bypassed the station.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today are to let the vacuum recover at TeV sector A1 and to begin its cryo system cooldown. Pelletron problems need to be resolved before putting the TeV into shot setup, but the plan is to establish a store Saturday evening. TeV experts reported that the A17 separators seemed okay.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/02/2007    email Fermilab